OpenStudy (anonymous):

A. –6 B. 6 C. 36 D. 72

OpenStudy (anonymous):

Which is the value of b solves the equation? \[b ^{-3}=\frac{ 1 }{ 216 }\]

OpenStudy (anonymous):

stop saying that your going to scare the crap out of me @Directrix

OpenStudy (michele_laino):

please note that: \[216=27*8=(3)^{3}*(2)^{3}=(3*2)^{3}=6^{3}\]

OpenStudy (anonymous):

@ispike plz help

OpenStudy (anonymous):

hey ghost do you talk?

OpenStudy (anonymous):

guess not

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

PLZ HELP I DON'T GET THIS CRAP

OpenStudy (michele_laino):

please note that I gave you an hint!

OpenStudy (anonymous):

OH THANKS FOR THE NOTE @Michele_Laino JUST FIGURED IT OUT

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks! @TeenWolfGirl

OpenStudy (anonymous):

np you helped out alot

OpenStudy (michele_laino):

ok! :)